### v1

# SAGe: A Lightweight Algorithm-Architecture Co-Design for Alleviating Data Preparation Overheads in Large-Scale Genome Analysis

Abstract.
Genome analysis plays an important role in many critical fields. Therefore, there have been extensive efforts to accelerate genome analysis by alleviating the challenges of analyzing exponentially growing volumes of genomic data. Prior works on accelerating genome analysis typically assume that the data is ready to be analyzed in the desired format; in real usage scenarios, however, it is common practice to store genomic data in storage systems in a compressed format. Unfortunately, preparing genomic data (i.e., accessing the compressed data from the storage system, decompressing the read data, and reformatting it) for a genome analysis accelerator leads to large performance and energy overheads and, thus, significantly diminishes the potential intended benefits of the acceleration efforts. To actually harness the full benefits of the extensive efforts on accelerating genome analysis, without resorting to storing massive volumes of genomic data uncompressed, there is a critical need to effectively address data preparation overheads. The solution must meet three criteria: (ihigh performance and energy efficiency for data preparation, (iihigh compression ratios, comparable to state-of-the-art genomic compression algorithms, and (iii) be lightweight to enable seamless integration with a broad range of genome analysis systems, regardless of where they are implemented. Simultaneously optimizing for all of these criteria is challenging, particularly due to the high decompression complexity of state-of-the-art genomics-specific compression algorithms (which, despite their high compression ratios, lead to large performance/energy overheads and require costly hardware resources) and resource constraints of a wide range of genome analysis systems.
We propose SAGe, an algorithm-system co-design that enables highly-compressed storage and high-performance access of large-scale genomic data in the desired format. Driven by our rigorous analysis of genomic datasets’ features, we propose an efficient and synergistic co-design of new (de)compression algorithm, lightweight hardware units, storage data layout, and interface commands. To achieve high performance, energy efficiency, and low hardware cost when preparing data, SAGe encodes data in lightweight structures that can be interpreted and decoded by efficient sequential scans and lightweight hardware. To maintain high compression ratios using only these lightweight structures, SAGe exploits unique features of genomic data to enhance space efficiency. SAGe’s lightweight design can be seamlessly integrated with a broad range of existing acceleration techniques for genome analysis to, for the first time, enable unleashing the performance and energy benefits of genome analysis acceleration, without relying on large-scale genomic data to be stored in uncompressed forms. Our evaluation shows that SAGe improves the average end-to-end performance and energy efficiency of state-of-the-art genome-analysis accelerators by 3.0$\times$–12.3$\times$ and 18.8$\times$–49.6$\times$, respectively, compared to when the accelerators rely on state-of-the-art decompression tools.
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; isbn: 978-X-XXXX-XXXX-X/XX/XX
1. Introduction
Genome analysis, which refers to the analysis of the genomic information encoded in an organism’s DNA, plays an important role in many critical fields, such as personalized medicine (clark2019diagnosis, ; farnaes2018rapid, ; sweeney2021rapid, ; alkan2009personalized, ; flores2013p4, ; ginsburg2009genomic, ; chin2011cancer, ; Ashley2016, ), tracing outbreaks of communicable diseases (bloom2021massively, ; yelagandula2021multiplexed, ; le2013selected, ; nikolayevskyy2016whole, ; qiu2015whole, ; gilchrist2015whole, ), and ensuring food safety through pathogen monitoring (e002244, ; TONG2021130, ). To be able to analyze genomic information computationally, a DNA sample from an organism undergoes a process called sequencing, which converts the information from DNA molecules in the sample to digital data. This digital information is encoded in the form of sequences that represent the DNA nucleotides with the characters A, C, G, and T. Current sequencing technologies cannot sequence long DNA molecules end-to-end. Instead, state-of-the-art sequencers generate randomly- and redundantly-sampled smaller and inexact fragments of these molecules, known as reads. Sets of genomic reads (called read sets) are then used in genome analysis. The importance of genome analysis, along with the rapid improvement in sequencing (i.e., reduced costs and increased throughput (berger2023navigating, )), has led to the rapidly increasing adoption of genome analysis in recent years (clark2019diagnosis, ; farnaes2018rapid, ; sweeney2021rapid, ; ginsburg2009genomic, ; chin2011cancer, ; Ashley2016, ; bloom2021massively, ; gilchrist2015whole, ) and continuous exponential growth in genomic data generation (katz2021sra, ; ena2022, ; srastats, ; enastats, ) (far outpacing Moore’s law (stephens2015big, )).
Due to its importance, there have been extensive efforts to accelerate genome analysis by alleviating the challenges of analyzing exponentially growing volumes of genomic data [^1] We focus on genomic read sets, which are the predominant form of genomic data (enastats, ).. These efforts focus on alleviating computational overheads via algorithmic optimizations (e.g., (zhang2000greedy, ; slater2005automated, ; li2018minimap2, ; myers1999fast, ; marco2021fast, ; marcosola2023optimal, ; grootkoerkamp2024apa2, )) or hardware accelerators (e.g., (mutlu2023accelerating, ; alser2022molecules, ; lou2020helix, ; lou2018brawl, ; shahroodi2023swordfish, ; saavedra2020mining, ; markus2020benchmarking, ; subramaniyan2021accelerated, ; huangfu2018radar, ; khatamifard2021genvom, ; cali2020genasm, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, ; cheng2018bitmapper2, ; houtgast2018hardware, ; houtgast2017efficient, ; zeni2020logan, ; ahmed2019gasal2, ; nishimura2017accelerating, ; de2016cudalign, ; liu2015gswabe, ; liu2013cudasw++, ; liu2009cudasw++, ; liu2010cudasw++, ; wilton2015arioc, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fujiki2020seedex, ; banerjee2018asap, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, )), and/or reducing data movement overheads via near-data processing (e.g., in main memory (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ) or storage system (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )).
Prior works on accelerating genome analysis typically assume that the data is ready to be analyzed in the desired format; in real usage scenarios, however, it is common practice to store genomic data in storage systems in a compressed format (berger2023navigating, ; zhu2013high, ; Deorowicz2013, ). This is because storing raw (i.e., uncompressed) data in multiple formats for different accelerators is impractical, given the massive volume of genomic data. In fact, due to the importance of efficiently storing large volumes of genomic data in a space-efficient manner, there exist many compression techniques (e.g., (chandak2018spring, ; chandak2017compression, ; kokot2022colord, ; Meng2023, ; roguski2018fastore, ; dufort2021renano, ; kowalski2019pgrc, ; dufort2020enano, )) specialized for genomic data to achieve higher compression ratios compared to state-of-the-art general-purpose compressors.
Unfortunately, preparing genomic data (i.e., accessing the compressed data from the storage system, decompressing the read data, and reformatting it) for a genome analysis accelerator significantly diminishes the potential intended benefits of genome analysis acceleration efforts. Our motivational analysis (§3) shows that this overhead can significantly increase the end-to-end execution time and energy consumption of genome analysis. In particular, we observe that the decompression overhead can be even more significant than the I/O overhead that occurs when reading larger uncompressed files from storage. Our observation highlights a critical trade-off; while it is possible to avoid decompression time by storing data in an uncompressed way, doing so requires allocating significantly larger storage capacity, which is inefficient, costly, and unsustainable for rapidly increasing volumes of genomic data.
To be able to actually harness the full benefits of the extensive efforts on accelerating genome analysis, without resorting to storing massive volumes of genomic data in uncompressed forms, there is a critical need for a design that effectively addresses data preparation overheads while meeting three key requirements. First, it should achieve high performance and energy efficiency when preparing data. Second, it should maintain high compression ratios comparable to state-of-the-art genomics-specific compression techniques. Third, it must be lightweight to enable high versatility for seamless and unobtrusive integration with a broad range of genome analysis systems, including general-purpose systems, specialized accelerators, and near-data processing architectures.
Simultaneously optimizing for all three requirements (i.e., high data preparation performance and energy efficiency, compression ratios comparable to state-of-the-art genomic compression, and lightweightness for seamless integration) is challenging. Some prior works (e.g., (guo2013gpu, ; qiao2019fpga, ; zhao2017streaming, ; jiang2021exma, ; arram2015fpga, ; wang2018accelerating, ; chen2023efficient, ; leavline2013hardware, )) accelerate certain computational kernels (e.g., BWT (zhao2017streaming, ; qiao2019fpga, ), FM-Index search (jiang2021exma, ; arram2015fpga, ; wang2018accelerating, ), and LZMA (chen2023efficient, ; leavline2013hardware, )) that are widely used in many state-of-the-art genomics-specific compression techniques (e.g., (lan2021genozip, ; Meng2023, ; kowalski2019pgrc, ; chandak2017compression, )). Despite their benefits, these approaches require expensive computational units, large buffers, or large DRAM bandwidth or capacity (e.g., due to many random accesses to large data structures for matching patterns). Therefore, none of these approaches is suitable for seamless integration with genome analysis accelerators, particularly when targeting integration with genome analysis systems in resource-constrained environments. For example, while near data processing (NDP) leads to significant benefits in genome analysis (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ; mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, ) by fundamentally alleviating data movement overheads when analyzing large-scale genomic data, such designs are often implemented in resource-constrained environments (e.g., within the memory system or storage device), which makes expensive decompression techniques prohibitive to deploy. Consider NDP accelerators implemented inside the storage system (e.g., (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )), which provide large performance and energy benefits for genome analysis by (i) fundamentally alleviating data-movement overhead from the storage system and (ii) reducing the application’s burden from the rest of the system (i.e., computational units and main memory). Without the capability to prepare data for analysis inside the SSD’s resource-constrained environment, all such works would either necessitate data to be stored in an uncompressed way, which is inefficient, or require external decompression, which completely undermines the fundamental benefits of NDP.
Our goal in this work is to improve the performance and energy efficiency of genome analysis by mitigating data preparation overheads, while maintaining compression ratios comparable to state-of-the-art genomics-specific compressors, and ensuring a lightweight design that can be seamlessly integrated with a broad range of genome analysis systems. To this end, we propose SAGe, an algorithm-system co-design that enables highly-compressed storage and high-performance access of large-scale genomic data in the desired format. The key idea of SAGe is to leverage the unique features of genomic data in both the compression mechanism and architecture design. SAGe’s lightweight co-design, for the first time, enables unleashing the performance and energy benefits of genome analysis acceleration, without relying on large-scale genomic data to be stored in uncompressed forms.
Key Mechanism. Driven by our rigorous analysis of genomic datasets’ features, we propose an efficient and synergistic co-design of new (de)compression algorithm, lightweight hardware units, storage data layout, and interface commands. SAGe’s co-design comprises four aspects. First, to achieve high performance, energy efficiency, and low hardware cost when preparing data, SAGe encodes data in lightweight structures that can be decoded by efficient sequential scans and lightweight hardware. To this end, we sequentially encode information regarding reads in a genomic read set in lightweight array structures. We efficiently leverage the long-range similarity common in DNA data to compactly, and yet losslessly, encode genomic reads in these structures. We reduce the size of the array structures by adapting the bit width of each entry. To enable correct decoding of data, we use a guide array that records the bit width used for each corresponding array entry. Second, to maintain high compression ratios using only these lightweight structures, SAGe exploits unique features of genomic data by tuning the encoding of array and guide array structures (e.g., bit count granularities) based on the characteristics of each genomic dataset, which are shaped by the sequencing technology and the dataset’s genetic composition. Third, we design lightweight hardware units, which can be seamlessly integrated with any genome analysis system to efficiently interpret and decode SAGe’s data structures with only simple operations and efficient streaming accesses. Fourth, we design an efficient data layout in the storage system that enables leveraging the SSD’s full bandwidth when accessing genomic data. Fifth, we propose specialized interface commands, which are exposed to genome analysis applications to communicate with SAGe’s hardware units to decompress and convert data to the accelerator’s desired format (e.g., two-bit (cheever1989using, ) or one-hot (Hirst1992, ) encoding).
We evaluate SAGe when preparing data for different genome analysis systems. We compare SAGe against configurations where data decompression is performed via (i) pigz (adler2015pigz, ), a widely-used general-purpose compression tool; (ii) Spring (chandak2018spring, ) and NanoSpring (Meng2023, ), state-of-the-art genomics-specific compression tools for short and long reads, respectively; and (iii) Spring (chandak2018spring, ) and NanoSpring (Meng2023, ) with hardware acceleration. We evaluate SAGe’s impact on the end-to-end performance and energy efficiency of genome analysis when integrated with a state-of-the-art accelerator for read mapping (chen2023gem, ) (a fundamental genome analysis task). We show that SAGe improves performance by an average of 12.3$\times$, 3.9$\times$, and 3.0$\times$, and energy efficiency by 49.6$\times$, 24.6$\times$, and 18.8$\times$ compared to configurations where pigz, (Nano)Spring, and hardware-accelerated (Nano)Spring prepare the data for the read mapping accelerator. We also demonstrate SAGe’s benefits in preparing data in resource-constrained environments and its capability for integration with genome analysis systems, regardless of where they are implemented. To this end, we evaluate SAGe’s integration with GenStore (mansouri2022genstore, ), a state-of-the-art NDP genomic accelerator inside the SSD. SAGe provides 30.7$\times$, 9.9$\times$, and 7.5$\times$ average speedup compared to the configurations with pigz, (Nano)Spring, and hardware-accelerated (Nano)Spring, respectively, with only the very low area cost of 0.7% of the three cores (cortexr4, ) in an SSD controller (samsung860pro, ). SAGe provides all of these benefits while achieving significantly larger compression ratios compared to pigz (2.9$\times$ larger, on average) and comparable compression ratios to the state-of-the-art genomics-specific compressors (average reduction of only 4.6%).
This work makes the following key contributions:
  • We demonstrate that preparing data (i.e., accessing data from the storage system and decompressing and reformatting it) for genome analysis accelerators can significantly diminish their end-to-end potential benefits.
  • We propose SAGe, the first algorithm-system co-design for high-compressed storage and high-performance access of large-scale genomic data, able to seamlessly integrate with a wide range of genome analysis systems. SAGe’s lightweight co-design, for the first time, enables unleashing the performance and energy benefits of genome analysis acceleration, without relying on large-scale genomic data to be stored in uncompressed forms.
  • We enable SAGe’s lightweight design by leveraging the features of genomic data in both compression mechanism and architecture design.
  • We rigorously evaluate SAGe and show that, when integrated with state-of-the-art genome sequence analysis systems, SAGe significantly improves their end-to-end performance and energy efficiency, while incurring only a low hardware cost.
2. Background
2.1. SSD Organization
Fig. 1 shows a modern NAND flash-based SSD’s organization.
Figure 1. Organizational overview of a modern SSD.
1
NAND Flash Memory. One or more NAND packages share a channel to communicate with the SSD controller. Each NAND package contains one or more (typically up to 4) flash dies or chips that share the NAND package’s I/O pins.
2
SSD Controller. The SSD controller has two key components. First, multiple embedded cores execute the flash translation layer (FTL). Second, per-channel flash controllers perform request handling and error correction for reliability.
3
DRAM. SSDs utilize an internal DRAM to improve SSD performance. It uses the DRAM to store (i) the logical-to-physical (L2P) mappings, (ii) recently-accessed data, and (iii) metadata for management tasks.
2.2. Genomic Workflow
Given a genomic sample, a typical workflow consists of three key steps (alser2020technology, ; alser2020accelerating, ; berger2023navigating, ): sequencing, basecalling, and analysis. The first step, sequencing, involves reading the DNA from a biological sample as digital signals. Due to limitations in current sequencing technologies, many organisms’ entire genomes cannot be read as complete, error-free sequences. Instead, these technologies produce many shorter sequences called reads that are randomly sampled from different locations in the genome, where the average number of reads per location is called the sequencing depth. Deeper sequencing provides more information per location, allowing later analysis steps to better distinguish between biological signals and sequencing errors. State-of-the-art sequencing technologies are distinguished by the characteristics of the reads they produce. Examples include Illumina sequencers that produce short (corresponding to 150 characters) reads with high accuracy ($\sim$99.9%) (glenn2011field, ; goodwin2016coming, ; quail2012tale, ; kchouk2017generations, ; pfeiffer2018systematic, ), PacBio HiFi sequencing (wenger2019accurate, ; hon2020highly, ; amarasinghe2020opportunities, ) and recent ONT sequencers (sereika2022oxford, ; ni2023benchmarking, ) that produce long (1k to 25k) reads with intermediate accuracy ($\sim$99%) (hon2020highly, ; ni2023benchmarking, ), and older ONT sequencers that produce long reads (1.5k to 7.5k (ni2023benchmarking, )) with lower accuracy ($\sim$85-94%) (wang2021nanopore, ).
The second step, basecalling, converts the raw digital signals from a sequencer into strings. The signal from each DNA read is converted into three strings (cock2009sanger, ): (i) a header containing identifiers for the sequencing experiment and the specific read, (ii) the DNA encoded using an alphabet representing the nucleotides (also called base pairs) A, C, G, T (alongside N to represent an uncertain position), and (iii) a quality score (ewing1998base, ) for each nucleotide, encoding its probability of being incorrect using ASCII characters. The strings from all reads in a sample are then written sequentially [^2] This format is called FASTQ (cock2009sanger, ), the most commonly used read set format (illuminafastq, ). and stored in a compressed form (§2.3) for future analysis steps.
The third step is genome analysis on a collection of read sets. Genomic read sets can be obtained by, for example, (i) sequencing the DNA in a sample and storing the reads into the SSD of a sequencing machine (gamaarachchi2022fast, ; minion21, ) or (ii) downloading the read sets to a local SSD from a repository (leinonen2010sequence, ; bokulich2020measuring, ). A typical analysis workflow quantifies significant sequence mismatches between the reads and a reference genome. This typically involves two key tasks: First, the computationally-expensive read mapping process finds the potential matching locations of reads in a reference genome. Second, variant calling uses the mapping information to determine which observed mismatches are likely to be true biological sequence variation (and not sequencing errors). [^3] Analysis workflows access all DNA bases from the reads, but only a small fraction of the corresponding quality scores. This is because read mapping, while using all bases, typically ignores quality scores (li2018minimap2, ; vasimuddin2019efficient, ). The subsequent variant calling step only needs quality scores from the positions surrounding mismatches (Poplin2018, ; li2008mapping, ).
In many cases, the analysis step for a single read set can be performed many times and at different times. Some applications (e.g., measuring the genetic diversity in a population (kostlbacher2021pangenomics, ; vandorp2020emergence, )) require analyzing the genetic differences between a read set belonging to an individual and many reference genomes. Other applications require repeating the analysis step many times (e.g., with different read mapping parameters (khayat2021hidden, ; berger2023navigating, ) or updated reference genomes (chen2024improved, ; aganezov2022complete, ; berger2023navigating, ; Siren2024, )) to improve accuracy. Due to the criticality of the analysis step and the challenges of analyzing large amounts of genomic data on conventional systems, a large body of works (e.g., (mutlu2023accelerating, ; alser2022molecules, ; lou2020helix, ; lou2018brawl, ; shahroodi2023swordfish, ; saavedra2020mining, ; markus2020benchmarking, ; subramaniyan2021accelerated, ; huangfu2018radar, ; khatamifard2021genvom, ; cali2020genasm, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, ; cheng2018bitmapper2, ; houtgast2018hardware, ; houtgast2017efficient, ; zeni2020logan, ; ahmed2019gasal2, ; nishimura2017accelerating, ; de2016cudalign, ; liu2015gswabe, ; liu2013cudasw++, ; liu2009cudasw++, ; liu2010cudasw++, ; wilton2015arioc, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fujiki2020seedex, ; banerjee2018asap, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, )) accelerate genome analysis tasks.
2.3. Storing and Accessing Genomic Data
Driven by continuing exponential drops in DNA sequencing costs (nhgri, ; pedro2021integration, ), genomic sequence data volumes are growing by an order of magnitude every few years, approaching and expected to exceed the data volumes generated by various social media platforms (stephens2015big, ; srastats, ; katz2021sra, ; enastats, ). Due to the importance of efficiently storing large volumes of genomic data in a space-efficient manner, there exist many compression techniques (e.g., (chandak2018spring, ; chandak2017compression, ; kokot2022colord, ; Meng2023, ; roguski2018fastore, ; dufort2021renano, ; kowalski2019pgrc, ; dufort2020enano, )) specialized for genomic data to achieve higher compression ratios (e.g., typically ranging from $\sim$2 to $\sim$40 (chandak2018spring, ; roguski2018fastore, ; dufort2020enano, ; dufort2021renano, )) than general-purpose compression techniques (e.g., typically ranging from $\sim$2 to $\sim$(chandak2018spring, )). These genomics-specific compression techniques leverage the typically high degree of data redundancy in genomic datasets (especially at high sequencing depths, as explained in §2.2) to maximize compression ratios.
Fig. 2 shows an example of a typical genomics-specific compression technique. To store each read set in a compressed way, genomics-specific compression tools typically represent the read set with a 1 consensus sequence and 2 the mismatches of each read in the read set compared to that consensus (chandak2017compression, ; dufort2021renano, ; kokot2022colord, ; kowalski2019pgrc, ). A consensus sequence is an approximation of the sequenced organism’s genome, and can either be a user-provided reference genome (dufort2021renano, ) or a de-duplicated string derived from the reads, representing the most likely character at each location (chandak2017compression, ; kowalski2019pgrc, ). [^4] Note that it is not sufficient to store only the consensus sequence to represent the read set because the consensus sequence alone does not capture the full variability in the original reads. Individual reads often contain unique differences, including sequencing errors (goodwin2016coming, ; wang2021nanopore, ) or biological variation (e.g., when multiple copies of the same chromosome from each parent are present in the DNA sample (Cheng2022, )). Given this consensus sequence, a lossless encoding of a read consists of 1 its matching position in the consensus sequence (many compression tools, e.g., (chandak2017compression, ; Meng2023, ; roguski2018fastore, ; kowalski2019pgrc, ), store matching positions in a sorted manner based on their order of appearance in the consensus sequence, which enables delta encoding to reduce storage requirements), 2 mismatch positions (delta-encoded), 3 mismatch bases and types (i.e., substitution, insertion, deletion), and 4 read length (relevant for long-read sets since they have variable read lengths). This encoding of the reads is better compressible using general-purpose compressors (chandak2017compression, ). State-of-the-art genomic compression techniques further compress the mismatch information using various general compression techniques.
Figure 2. Overview of genomics-specific compression.
Note that identifying matching positions against the consensus sequence while compressing a read set does not eliminate the need for read mapping during subsequent analysis stages (Siren2024, ; Sedlazeck2018, ; Abel2020, ). This is due to three reasons. First, when compressing and storing a sequenced and basecalled read set, the exact reference genomes required for potential future read mapping analyses may not be known in advance (Siren2024, ). Second, a read set may undergo read mapping multiple times or at different times (as explained in §2.2). Therefore, simply having the matches against a single consensus sequence during compression is insufficient. Third, during compression, only one matching position for each read is stored (to minimize data size), whereas various read mapping use cases require many matching positions of each read (Sedlazeck2018, ; Abel2020, ; li2018minimap2, ).
3. Motivation
We perform experimental studies to understand the impact of preparing genomic data (i.e., accessing the data from storage and decompressing and reformatting it) for a genome analysis accelerator on end-to-end performance.
3.1. Methodology
Genome Analysis Accelerator. We use a state-of-the-art hardware accelerator (chen2023gem, ) for read mapping. To find the read mapping throughput, we use the throughput reported by the original paper.
Data Preparation. We use Spring (chandak2018spring, ) as the state-of-the-art genomic compression tool, and pigz (adler2015pigz, ) (parallel gzip (gailly1992gnu, )) as a widely-used general-purpose compression tool. For our evaluated real-world human read set (§6), [^5] The read set is sequenced from the widely-used 1000 Genomes Project benchmark dataset (10002015global, ) of representative human samples. Given the high degree of genetic similarity between human genomes (geneticsfacts, ), the evaluated human read set reflects the properties and trends commonly observed in most human populations. Spring achieves a 40.2$\times$ compression ratio, while pigz achieves 12.5$\times$.
To find the performance of data access from the SSD and decompression, we use a state-of-the-art SSD (samsungPM1735, ) with PCIe Gen4 (PCIE4, ) and a high-end server with AMD EPYC 7742 CPU (amdepyc, ) with 128 physical cores and 1.5-TB DDR4 DRAM. We use the best-performing thread count (i.e., the thread count after which adding more threads does not improve performance) for Spring (chandak2018spring, ) and pigz (adler2015pigz, ).
End-to-End Performance. We find the end-to-end throughput of the genome analysis task based on the throughput of preparing data for the accelerator and the throughput of the read mapping accelerator itself. I/O operations, decompression and reformatting, and read mapping operate in a pipelined manner and in batches (i.e., when decompressing and reformating reads in $batch\#i$, the mapper analyzes $batch\#i-1$), which enables partially overlapping parts of these steps. Further methodology details are in §6.
Evaluated Configurations. To show the impact of data preparation, we analyze configurations with different initial states for the input read set, each incurring different decompression and I/O overheads: (i) Cmprs1+I/O and (ii) Cmprs2+I/O with data stored in pigz- and Spring-compressed formats, respectively, within the storage system. (iii) Cmprs1+NoI/O and (iv) Cmprs2+NoI/O with data stored in pigz- and Spring-compressed formats, respectively, but in an idealized scenario with zero performance overhead due to storage I/O; (v) NoCmprs+I/O with data stored in an uncompressed and desired format for the accelerator in the storage system; and (vi) NoCmprs+NoI/O with data stored in an uncompressed format and zero performance overhead caused by storage I/O. This configuration assumes that the data is decompressed, in the desired format, and ready to be fed to the accelerator. We show each system’s throughput normalized to the throughput of NoCmprs+NoI/O.
3.2. Observations
Figure 3. Read mapping throughput with six configurations for a read set’s initial state.
Fig. 3 shows end-to-end throughput of preparing data and performing read mapping on it. We make four key observations. First, the overhead of preparing data for the accelerator significantly hurts end-to-end performance. Cmprs1+I/O and Cmprs2+I/O, which suffer from throughput loss due to I/O and decompression overheads, lead to 51.5$\times$ and 27.0$\times$ slowdown compared to NoCmprs+NoI/O, respectively. Second, the overhead of decompressing data to the desired format is larger than the I/O overhead. Removing the I/O overhead (in the Cmprs1+NoI/O and Cmprs2+NoI/O configurations) does not improve performance compared to Cmprs1+I/O and Cmprs2+I/O, as the decompression step dominates the end-to-end throughput of the data preparation pipeline. Third, decompression overhead has a larger impact on end-to-end throughput than the I/O overhead of reading larger uncompressed data from the storage system. Although NoCmprs+I/O requires reading larger uncompressed data from storage and thus incurs larger I/O overhead than Cmprs1+I/O and Cmprs2+I/O, it still achieves higher throughput than them (by 20.3$\times$ and 10.6$\times$, respectively). However, it is inefficient and unsustainable to store large volumes of genomic data uncompressed in the storage system. Fourth, I/O overhead, although less severe than the overhead of decompressing data into the desired format, still considerably affects end-to-end performance. NoCmprs+I/O leads to 2.5$\times$ slowdown compared to NoCmprs+NoI/O. Note that this I/O overhead is independent of the relative size of the read set and main memory since this I/O overhead is due to the need to move large amounts of low reuse data from the storage system to the read mapping accelerator (mansouri2022genstore, ). As shown in Fig. 3, the I/O overhead exists even when the main memory is larger than the read set.
3.3. Our Goal
Based on our observations, we conclude that data preparation leads to large overheads in the end-to-end performance of genome analysis, and can significantly diminish the potential benefits of genome analysis accelerators. Therefore, to be able to actually harness the full benefits of the extensive efforts on accelerating genome analysis, there is a critical need for a design that effectively addresses data preparation overheads while meeting three key requirements: (ihigh performance and energy efficiency for data preparation, (iihigh compression ratios, comparable to state-of-the-art genomic compression algorithms, and (iii) be lightweight to enable seamless integration with a broad range of genome analysis systems, regardless of where they are implemented.
Unfortunately, it is challenging to optimize both the performance of preparing data and the space-efficiency of storing data while maintaining low hardware costs. This is particularly challenging when targeting integration with genome analysis systems in resource-constrained environments. For example, while near data processing (NDP) leads to significant benefits in genome analysis (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ; mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, ) by fundamentally alleviating data movement overheads when analyzing large-scale genomic data, such designs are often implemented in resource-constrained environments (e.g., within the memory system or storage device), which makes expensive decompression techniques prohibitive to deploy. Consider NDP accelerators implemented inside the storage system (e.g., (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )), which provide large performance and energy benefits for genome analysis by (i) fundamentally alleviating data-movement overhead from the storage system (and thus help alleviate the I/O overhead in data preparation), and (ii) reducing the application’s burden from the rest of the system (i.e., computational units and main memory). Without the capability to prepare data for analysis inside the SSD’s resource-constrained environment, all such works would either necessitate data to be stored in an uncompressed way, which is inefficient, or require external decompression, which completely undermines the fundamental benefits of NDP.
Some prior works (e.g., (guo2013gpu, ; zhao2017streaming, ; jiang2021exma, ; arram2015fpga, ; chen2023efficient, ; leavline2013hardware, ; qiao2019fpga, ; wang2018accelerating, )) accelerate certain computational kernels, such as BWT (qiao2019fpga, ; guo2013gpu, ; zhao2017streaming, ), FM-index search (jiang2021exma, ; arram2015fpga, ), or LZMA (chen2023efficient, ; leavline2013hardware, ), which are widely used in many state-of-the-art genomics-specific compression tools (e.g., (lan2021genozip, ; Meng2023, ; kowalski2019pgrc, ; chandak2017compression, )). Despite their benefits, these works have two limitations. First, existing techniques mainly rely on many random accesses to large data structures (due to random look-ups for matching patterns in these structures), which makes their adoption inefficient in resource-constrained environments. For example, some works (e.g., (zhao2017streaming, ; wang2018accelerating, ; dragenora, )) demand high DRAM bandwidth. Some works require expensive hardware components, such as large buffers (qiao2019fpga, ), DRAM capacity (jiang2021exma, ), or large computation resources (arram2015fpga, ; guo2013gpu, ; leavline2013hardware, ). Therefore, these techniques are prohibitive to support in resource-constrained environments, such as portable genome analysis devices or NDP accelerators. [^6] For example, an NDP accelerator inside the storage system is implemented inside the resource-constrained environment of an SSD, with the limited-bandwidth single-channel internal DRAM (zou2022assasin, ) with small capacity (e.f., 1-GB DRAM for each TB capacity. For example, a 4-GB LPDDR4 DRAM is used in a modern 4-TB SSD (samsung860pro, ). While one can reduce the required buffer/DRAM spaces by decreasing window sizes, such an approach consequently lowers compression ratios (qiao2019fpga, ), which is not acceptable for adoption for genomic data. Second, in addition to their adoption limitations, these tools do not fully alleviate the end-to-end overhead of data preparation (as shown in §7) since they do not consider all of the steps involved in data preparation.
Our goal in this work is to improve the performance and energy efficiency of genome analysis by mitigating data preparation overheads, while maintaining compression ratios comparable to state-of-the-art genomics-specific compressors, and ensuring a lightweight design that can be seamlessly integrated with a broad range of genome analysis systems, such as general-purpose systems, specialized accelerators (mutlu2023accelerating, ; alser2022molecules, ; lou2020helix, ; lou2018brawl, ; shahroodi2023swordfish, ; saavedra2020mining, ; markus2020benchmarking, ; subramaniyan2021accelerated, ; huangfu2018radar, ; khatamifard2021genvom, ; cali2020genasm, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, ; cheng2018bitmapper2, ; houtgast2018hardware, ; houtgast2017efficient, ; zeni2020logan, ; ahmed2019gasal2, ; nishimura2017accelerating, ; de2016cudalign, ; liu2015gswabe, ; liu2013cudasw++, ; liu2009cudasw++, ; liu2010cudasw++, ; wilton2015arioc, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fujiki2020seedex, ; banerjee2018asap, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, ), and near-data processing architectures (e.g.,in main memory (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ) or storage system (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )).
4. SAGe: Overview
We propose SAGe, a scheme for storing and acessing large-scale genomic data that not only stores genomic data in a space-efficient compressed way, but also serves requested data in the desired format in a fast and energy-efficient manner.
We address the challenges of optimizing both the performance of preparing data and the space-efficiency of storing data while maintaining low hardware costs via holistic co-design of new (de)compression algorithm, light- weight hardware units, storage data layout, and interface commands, leveraging the features of genomic data. Doing so enables rapid interpretation and decoding via lightweight hardware, while relying on efficient sequential accesses when accessing stored genomic data. [^7] Since compression time is not on the critical-path of genome analysis pipelines, SAGe’s compression and encoding is performed on the host system. When writing the SAGe-compressed data, SAGe leverages its customized interface commands (§5.3) and FTL (§5.4) to efficiently map the data for its efficient access during decompression.
Overview. SAGe supports data preparation for genome analysis systems, regardless of where they are implemented. For example, Fig. 4 shows SAGe’s implementation inside and outside an SSD. The in-storage implementation of SAGe enables in-storage processing for genome analysis in practice (i.e., without relying on large-scale genomic data to be stored in uncompressed formats). On the other hand, SAGe’s implementation outside the SSD can provide benefits when two conditions are met: (i) the given genomic dataset and application do not benefit from ISP, and (ii) the SSD’s external bandwidth limits application performance (e.g., when an in-storage filter cannot efficiently work for the target dataset, and the bandwidth between the SSD and the genome analysis accelerator outside the SSD is limited). In such a case, SAGe’s lightweight hardware can be tightly integrated with the genome analysis accelerator to quickly prepare data and pass it to the accelerator at a high bandwidth.
Fig. 4(a) shows a high-level overview of SAGe’s operations, in the case of its in-storage data preparation. When the host requests genomic data ( 1 in Fig. 4(a)) using SAGe’s interface commands (§5.3), SAGe prepares for its operations based on the SAGe FTL (§5.4) metadata ( 2). The SAGe FTL manages communication between the SSD and the genome analysis systems and coordinates data flow between different SSD parts (e.g., flash chips and in-storage accelerators). After preparation, SAGe accesses genomic data in its SAGe-compressed format (§5.1) from flash chips ( 3) and decompresses it into the desired format using SAGe’s accelerators (§5.2) in the SSD controller ( 4). Based on the interface commands, the decompressed data can then go through genome analysis inside ( 5) and/or outside the SSD ( 6).
Figure 4. Overview of SAGe with its accelerator unit (a) inside and (b) outside the SSD.
Based on our design, SAGe’s data preparation outside the SSD is straightforward: As shown in Fig. 4(b), SAGe’s accelerators need to be implemented outside the SSD and use SAGe-specific I/O interfaces and FTL to manage data layout directly inside the SSD. Due to its lightweight design, SAGe’s hardware units outside the SSD can easily be integrated with the genome analysis systems either directly on the same chip (i.e., with the genome analysis accelerator, which can be a stand-alone ASIC (turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, ), or implemented in an FPGA (saavedra2020mining, ; zhang2023genomix, ; cervi2022metagenomic, ; shih_haru_2023, ; fujiki2020seedex, ; banerjee2018asap, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, ; liyanage2023efficient, ), near memory (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ), etc.) or connected via a PCIe interface (i.e., similar to how individual accelerators or GPUs connect to the system).  [^8] In our evaluations, we consider the PCIe-connected configuration for when SAGe’s hardware units are outside the SSD.
For its implementation in or near storage, it is possible for SAGe’s operations to run either on our lightweight specialized accelerators or, alternatively, on other general-purpose in-storage (e.g., (torabzadehkashi2019catalina, ; zou2022assasin, )) or near-storage (e.g., Samsung SmartSSD (lee2020smartssd, )) processing systems. This is because, leveraging our optimizations in algorithms and data structures, SAGe’s operations require only simple operations and small buffers. For example, SAGe’s hardware units consume only 2.5% of the lookup tables and 0.8% of the flip-flop resources of SmartSSD’s FPGA (lee2020smartssd, ). [^9] Since SmartSSD’s FPGA is connected to the SSD with a PCIe lane, it can act as the SGout (SAGe outside the SSD) configuration analyzed in Fig. 12 and provide similar benefits. This emphasizes SAGe’s ease of adoption. Ultimately, choosing between these SAGe configurations (specialized lightweight hardware or general in-/near-storage units) is a design decision with different trade-offs, with the accelerator achieving higher power efficiency. Efficiently leveraging these underlying hardware units relies on SAGe’s optimized data mapping, interface commands, and algorithms.
5. SAGe: Detailed Design
5.1. Compression and Data Structures
Figure 5. Overview of SAGe’s compression mechanism.
Fig. 5 shows an overview of SAGe’s compression mechanism. SAGe also takes the consensus-based approach widely used in existing genomics-specific compression techniques (chandak2017compression, ; dufort2021renano, ; kokot2022colord, ; kowalski2019pgrc, ); for each read, SAGe stores the mismatch information between the read and a consensus sequence ( 1 in Fig. 5). [^10] Note that, as detailed in §2.3, identifying matches against the consensus during compression does not eliminate need for read mapping during later analyses (Siren2024, ; Sedlazeck2018, ). The key difference of SAGe over existing techniques is how it encodes the mismatch information; instead of using expensive general-purpose compression algorithms (e.g., LZMA (pavlov2015lzma, ), the BWT-based BSC (grebnov2011libbsc, ) or FM-Index search (ferragina2000opportunistic, ), etc.), SAGe (i) stores the information in simple structures that can be efficiently decoded by lightweight hardware with streaming data access, and (ii) minimizes the sizes of these structures by leveraging the features of read sets.
To this end, SAGe sequentially stores the mismatch information (in this figure’s example, delta-encoded mismatch positions) of different reads back-to-back ( 2) in a mismatch position array (MPA). SAGe then improves the capacity efficiency ( 3) by (i) picking the most suitable set of bit count values to use for each entry in the MPA and (ii) using a lightweight structure, i.e., mismatch position guide array (MPGA), to encode which bit count has been used to store each entry in the MPA. In this example, a value of 0 or 1 at index $i$ of the MPGA specifies that the corresponding entry in the MPA is stored using one or four bits, respectively. This is based on the key observation that the distribution of delta-encoded values of mismatch positions is heavily skewed towards smaller values (reasons in §5.1.1). During compression, to minimize the overall data size, SAGe tunes the parameters of each array and guide array ( 4) based on the features of each read set, such as read length and mismatch distribution, which depend on the sequencing technology and the genetic composition of the read set. The configuration parameters of the array and guide array bit count values used for compressing each read set are then encoded at the beginning of the corresponding compressed read set file. When the read set is accessed, SAGe decodes these parameters and utilizes them during the decompression process (§5.2.3).
This data encoding enables decoding with only efficient sequential scans through the data structures. Since SAGe stores mismatch information in the order of appearance in each read, decoding this information is feasible via a sequential scan through the array structures. Since SAGe stores the reads’ positions based on their order of appearance in the consensus sequence, [^11] As discussed in §2.3, various other genomic compressors (e.g., (chandak2017compression, ; chandak2018spring, ; roguski2018fastore, ; cogo2021genodedup, )) also store reads in this order, which enables delta encoding for capacity-efficiency. SAGe also only needs sequential scans through the consensus. Due to this efficient encoding, SAGe can directly send each read to the genome analysis system as soon as it is decoded. Since genome analysis applications can analyze reads independently, they can begin analyzing reads as soon as they are received from SAGe.
By leveraging all of the features that enable SAGe to decode data using only lightweight operations and efficient scans through the data, SAGe is also inherently memory-hierarchy friendly on general-purpose systems due to two reasons. First, as opposed to the state-of-the-art compression techniques used for genomic data that require a large number of random accesses to the main memory, SAGe relies only on efficient streaming accesses. Second, given that all data is accessed sequentially with full spatial locality, each cache line is fully used, which leads to fewer accesses to main memory compared to other algorithms that require many fine-grained random accesses. These factors make SAGe suitable also for cases where the user wants to use it purely in software on general-purpose systems.
5.1.1. Mismatch Positions and Mismatch Counts.
SAGe stores mismatch positions for different reads in a read set sequentially in MPA and uses MPGA to indicate the number of bits used for each MPA entry. SAGe stores the number of mismatches in each read followed by the mismatch positions in that read. This allows SAGe to determine how many elements of MPA and MPGA it needs to access in order to retrieve all mismatch positions for each read. To reduce the sizes of MPA and MPGA, SAGe introduces data-driven optimizations that leverage the distinctive characteristics of short and long read sets. Fig. 6 demonstrates these characteristics for representative real-world genomic read sets (details in §6).
Figure 6. (a) Distribution of #bits needed to store the delta-encoded mismatch positions, (b) Distribution of #reads with different mismatch counts, (c) Cumulative distribution of indel block lengths, (d) Cumulative distribution of #bases in indel blocks of different lengths.
First, Fig. 6(a) shows the distribution of the number of bits required to store the delta-encoded mismatch positions for a representative long read set (RS4 in Table 3). We observe that the distribution is skewed towards the smaller values. This is due to two main factors that can lead to nearby mismatches. First, mutations tend to cluster in some regions of the genome (Bourque2018, ). Second, a local drop in base quality during sequencing can cause incorrect bases in the read (LaPierre2019, ). Based on this observation, SAGe performs two optimizations to minimize the overall size of MPA and MPGA. First, during compression for each read set, SAGe tunes (i) the number of distinct bit counts used for MPA, and (ii) the specific values of each bit count. To do so, SAGe generates the distribution of bits needed to store all mismatch positions in a read set. Based on this distribution and examining various combinations of the bit count values used in MPA and MPGA, SAGe selects the combination that minimizes the total size of MPA and MPGA in the read set. Second, SAGe refines the guide array encoding by exploiting the higher frequency of smaller bit counts. For example, with four different bit counts ($N_{0}$ to $N_{3}$), rather than using 00, 01, 10, and 11 to represent them, SAGe uses 0, 10, 110, and 1110. This encoding can then be efficiently decoded via sequentially scanning MPGA.
Second, we analyze the mismatch counts per read in Fig. 6(b) for a representative human short read set (RS2 in Table 3). As also shown by prior works (e.g., (nag2019gencache, ; mansouri2022genstore, )), most short reads have no or few mismatches. To alleviate the overhead of storing #mismatches for most reads, we use the adaptive encoding mentioned above.
Third, recognizing that indel blocks (consecutive inserted or deleted bases) cause many mismatches in long reads, we analyze their characteristics in a representative read set (RS4 in Table 3). Figures Fig. 6(c) and Fig. 6(d) show the cumulative distributions of indel block lengths and the bases stored for different block lengths, respectively. We observe that (i) most indel blocks are of length one—a common trait in read sets (ono2020pbsim2, ; belyaeva2022best, ), and (ii) although single-base blocks are more frequent, larger blocks encompass most indel bases. Based on these, we implement two optimizations. First, we store the position of the first mismatch and the length of the indel, rather than storing each mismatch position. Second, after detecting an indel (see §5.1.2), we reserve one bit in MPGA to indicate if it is a single-base indel; if not, we allocate 8 bits in MPA for the block length, thus saving space for most single-base indels.
Fig. 7 shows an example of the mismatch position data structures and decompression process for long reads, demonstrating MPGA ( 1), MPA ( 2), and how MPGA entries encode bit counts in MPA ( 3). To retrieve mismatch positions, SAGe first reads the mismatch count for Read#1 in MPGA. It then scans MPGA to obtain the bit counts for each mismatch position and reads the specified number of bits from MPA to decode these positions. For example, if MPGA’s entry for a mismatch position is 10, SAGe reads 4 bits in MPA to find that mismatch position. If a mismatch is an indel (e.g., Read#1 Mismatch#2), SAGe checks MPGA if the indel length is greater than 1. If so, it reads the next eight bits from MPA to determine the indel length. After decoding all three mismatch positions of Read#1, SAGe proceeds to the next read. These lightweight structures enable SAGe to store mismatch positions in a capacity-efficient way and decode them using sequential scans.
Figure 7. Example of the data layout and decoding of mismatch positions for long reads in SAGe.
5.1.2. Mismatching Bases and Types.
SAGe stores information regarding mismatching bases/types in a mismatching base and type array (MBTA). Similar to MPA, MBTA is filled sequentially with information on mismatches in different reads. SAGe includes two optimizations in MBTA . First, in long reads, we observe that a significant fraction of mismatching bases (e.g., up to 80% in our datasets) can originate from chimeric reads, which are reads with sequences joined from different regions of the genome due to sequencing or library preparation errors or structural variations (guan2016structural, ). Parts of these reads map to different locations in the consensus sequence. Fig. 8 shows an example of a chimeric read and its two matching positions. The read has eight mismatches in matching position#1 and nine mismatches in matching position#2. Therefore, only accounting for the best matching position (i.e., matching position#1 in this example), as commonly done in prior works (e.g., (dufort2021renano, ; kowalski2019pgrc, )), results in a large number of mismatches. To avoid this overhead, SAGe considers the top $N$ matching positions [^12] In our analysis, $N=3$ leads to the best compression ratio. This parameter can be tuned seamlessly. for chimeric reads and reconstructs the read based on them. For example, considering both matching positions in Fig. 8 leads to only three mismatches, which is smaller than the eight mismatches found at the best matching position. Since, in real chimeric long reads, the number of mismatches at individual matching positions can be substantial, it is more efficient to reconstruct the read from several matching positions (and store the extra matching positions), rather than storing a large number of mismatch bases at only the top matching position.
Figure 8. Example of a chimeric read.
Second, since in short reads, substitutions constitute a large fraction of mismatches, we devise an optimization that eliminates the need to store types for substitutions. Instead of using two bits to encode each base (A, C, G, T) and two bits to encode each type (substitution, insertion, deletion), we merge base and type encodings to avoid storing types for substitutions. When reconstructing a read during decompression, SAGe checks whether the mismatching base is the same as the consensus sequence at the corresponding mismatch position. If it is, SAGe recognizes that the mismatch is not a substitution but an indel and uses one bit after the mismatching base to distinguish between an insertion or a deletion.
5.1.3. Matching Positions.
Fig. 9 shows the distribution of the number of bits needed to store delta-encoded matching positions for a representative real human short read set. We observe that most delta-encoded matching positions can be encoded with very few bits. This is due to its sequencing depth, which results in genomic reads being closely spaced, leading to small differences between consecutive matching positions. Since mismatch positions also exhibit a similar skew towards smaller values (as shown in §5.1.1), we apply the same optimization used for mismatch positions to matching positions as well by introducing the matching position array (MaPA) and matching position guide array (MaPGA).
Figure 9. Distribution of the number of bits required to store the delta-encoded matching position.
5.1.4. Corner Cases.
Although most reads are compressed using the encodings in §5.1.15.1.3, two mismatch types require special handling. First, reads containing unidentified bases (N) expand the DNA alphabet to five characters, making 2-bit encoding impossible. Second, some reads include clips—large insertion blocks at the beginning or end. We classify a read as a corner case if it has a mismatch at position 0. We then use a single bit in MBTA to indicate whether a read has a mismatch at position 0 or is a corner case.
5.1.5. Quality Scores.
Similar to various other genomic compressors (e.g., (chandak2017compression, ; Meng2023, )), SAGe does not include specialized support for quality scores and, instead, can integrate with any quality score (de)compressor (e.g., (dufort2020enano, ; kokot2022colord, )) outside the SSD. This is because, despite the importance of compression ratios for quality scores, quality score decompression time is not on the critical path of many genome analysis pipelines.
5.2. SAGe’s Hardware Design
SAGe’s accelerator recognizes SAGe’s data structures and quickly decompresses them to provide the application with data in its desired format. Fig. 10 shows SAGe’s hardware units and their integration with the SSD in the case of their in-SSD implementation. We discuss SAGe’s data mapping and fetching from the SSD (§5.2.1), the hardware components’ details (§5.2.2), and their operations (§5.2.3). For its outside-SSD implementation, SAGe places the hardware components outside the SSD but still relies on its efficient data mapping inside the SSD.
5.2.1. Mapping and Fetching Data.
SAGe optimizes data mapping in the SSD and data fetching to the hardware accelerator such that (i) the accelerator leverages SSD’s full bandwidth and (ii) does not require costly hardware resources (e.g., large buffer spaces).
To efficiently leverage all SSD channels, SAGe uniformly partitions the data between the flash channels. This efficient partitioning is enabled by SAGe’s sequential access pattern to all of its data structures. Each partition of the consensus sequence, along with the compressed mismatch information of the reads mapped to that partition, is then placed in a separate channel.
To enable accessing data without requiring large buffer spaces, which is particularly critical for when SAGe’s hardware units need to integrate with genome analysis systems in resource-constraint environments (e.g., near-data processing systems within the SSD), SAGe adopts two optimizations. First, we adopt an approach similar to (zou2022assasin, ) and operate on data fetched from NAND flash chips without needing to buffer them in the SSD’s low-bandwidth internal DRAM. Second, as we no longer rely on DRAM buffers, we must minimize the SRAM buffer sizes necessary for performing computations directly on NAND flash data streams. We do so based on two key features enabled by SAGe’s algorithm and data structures: (i) SAGe’s operations are lightweight and perform only streaming accesses, thus do not need to buffer large amounts of data during computation, and (ii) SAGe employs per-channel accelerator units, and since data can be uniformly partitioned across channels, these units operate independently without requiring synchronization or inter-channel buffering. Leveraging these features, SAGe performs operations on flash data streams using a lightweight double-buffering technique. We use one register to load a chunk of data, and another to hold the subsequent data chunk being read from the flash chips. For this double-buffering, we use two 64-bit registers as they are sufficient to fully utilize our per-channel accelerators.
To leverage the full bandwidth of each channel, We devise a specialized data mapping scheme. To this end, we store different SAGe data structures (i.e., the consensus sequence, guide arrays, and position arrays) sequentially in an interleaved manner, such that SAGe only requires sequential access through data in each channel. When a chunk of data is ready in the 64-bit register, SAGe’s hardware accelerator distributes the arrays and guides arrays to the respective registers in the accelerator’s components (§5.2.2) based on the values indicated by the guide arrays (as explained in §5.1). Details on SAGe FTL’s logical-to-physical mapping in flash blocks across different channels are provided in §5.4.
Figure 10. Overview of SAGe’s hardware units.
5.2.2. Components.
Fig. 10 shows the structure of SAGe’s hardware accelerator, consisting of three key components. 1 Scan Unit (SU) sequentially scans through position arrays to find the mismatch information. To do so, it scans through the position guide array and as soon as it reaches a zero bit, it detects that a new element of the guide array is read. Based on this element’s value, SU determines how many bits to read from the position array (e.g., as shown in the example of Fig. 7). The SU then adds the newly-read value of the position array to the previously decoded mismatch position to find the new mismatch position. 2 Read Construction Unit (RCU) receives the position information from SU and reconstructs the full reads by plugging the mismatches into the correct positions of the consensus sequence. RCU sequentially scans through the consensus sequence and whenever it reaches a mismatch position, it replaces the base in that position with the mismatching base. 3 Control Unit (CU) coordinates operations between SU and RCU.
Due to the streaming access patterns of SU and RCU operations, they only require small registers to buffer a small part of each array. The Guide Array Register and Array Registers each consist of eight bits since that is the largest element size used in the arrays in our compression scheme design. The Configuration Register is also eight bits since we load the configuration parameters in eight-bit chunks. We consider a Consensus Sequence Register of size 150 base pairs since this is the largest read size in most short read sequencing datasets (glenn2011field, ; quail2012tale, ). For short reads longer than this size and for long reads, we reconstruct the reads in 150-base-pair chunks.
5.2.3. Operations.
When receiving a request to access a read set, SAGe’s per-channel accelerators start streaming data from flash channels. First, SU reads the configuration parameters of the arrays and guide arrays ( 1 in Fig. 10). This includes the encoded associations between arrays and guide arrays (e.g., between MPGA and MPA, as shown in Fig. 7). Second, SU scans through the position arrays ( 2) as detailed in §5.2, retrieves the matching positions and mismatch positions ( 3), and sends them to RCU ( 4). Meanwhile, RCU scans through the consensus sequence and the base/type array ( 5). Whenever RCU detects an indel (i.e., when the mismatch base in MBTA equals the consensus sequence base at the corresponding mismatch position, as detailed in §5.1.2), it sends a signal to SU to decode information about the indel and its length accordingly 6. Finally, based on all the mismatch information (i.e., mismatch position, type, and base), RCU reconstructs each read by substituting or inserting mismatches into the correct positions in the consensus sequence, or by deleting the relevant bases if the mismatch is a deletion ( 7). Based on the interface command (§5.3) from the genome analysis system, RCU flexibly constructs the reads in the desired format and sends them for genome analysis ( 8).
5.3. Interface Commands
SAGe needs three new NVMe commands.
SAGe_Read: It is a specialized read command that (i) is used to access genomic data, and (ii) specifies the output format (e.g., 2-bit or 1-hot encoding). The other parameters needed for SAGe (e.g., array parameters) are written at the beginning of each compressed file and are loaded into SU (§5.2) when accessing a file. Upon receiving this command, SAGe uses SAGe FTL (§5.4) for its operations.
SAGe_Write: It is a specialized write command that is used when writing genomic data to the SSD and updates SAGe FTL’s mapping metadata.
SAGe_ISP: It specifies whether the data, after being prepared by SAGe, should go to genomic ISP units or whether it should directly go outside the SSD for further analysis.
5.4. SAGe’s Flash Translation Layer
SAGe FTL needs simple changes to the conventional FTL for handling genomic data and does not interfere with the baseline vendor-specific features of the FTL when handling non-genomic data. SAGe FTL dedicates each block in the SSD to either genomic data or non-genomic data. The SSD detects genomic data is being accessed when it receives a request via SAGe’s interface commands (§5.3). For all other types of data, SAGe does not impact the functionality of the baseline FTL and controller, and the SSD can be accessed similarly to a conventional general-purpose SSD.
Data Placement. SAGe performs data placement for genomic data to leverage the maximum throughput of the SSD and the accelerators of SAGe. To this end, when writing a compressed genomic read set, SAGe uniformly partitions the data across the flash channels in a round-robin manner. This efficient partitioning is enabled by SAGe’s sequential access pattern to all of its data structures. Each partition of the consensus sequence, along with the compressed mismatch information of the reads mapped to that partition, is then placed in a separate channel. Fig. 11 shows an example of how compressed genomic datasets are organized in blocks across SSD channels, dies, and planes. When storing a compressed read set across all channels, SAGe FTL writes data in a round-robin manner such that the active block (i.e., blocks available for writes) in different channels have the same page offset. This placement then enables SAGe FTL to simply perform efficient multi-plane read operations when accessing data across all channels and leverage the SSD’s full bandwidth. With this efficient data placement, and given that each genomic read in a genomic dataset can be analyzed independently, SAGe concurrently streams through data from each channel and performs decompression and reformatting (§5.2).
Figure 11. Data layout in SAGe.
SSD Management Tasks. When accessing and preparing genomic data, SAGe does not write any new data and only accesses the read set sequentially. Therefore, during this phase, SAGe does not need to perform write-related management (e.g., garbage collection (kim2020evanesco, ; cai2017error, ; park-dac-2016, ; park-dac-2019, ) and wear-leveling (chang2007efficient, ; cai-insidessd-2018, )). SAGe’s accelerators are located in the SSD controller and access data after ECC decoding.
When performing garbage collection on genomic data, we must select victim blocks such that we preserve SAGe’s capabilities for efficient multi-plane operations. This requires maintaining the same page offset across all blocks in a parallel unit. Given that genomic read sets do not require partial updates, and that they are accessed sequentially, it is straightforward to realize this efficient garbage collection. SAGe FTL performs garbage collection in a grouped manner and selects every block in the parallel unit as a group of victim blocks, which are then sequentially rewritten. SAGe selects this group of victim blocks in the order they were originally written, as indicated by their logical address sequence in the dataset.
SAGe performs other management tasks needed for ensuring reliability (e.g., refresh to prevent uncorrectable errors (luo2018improving, ; luo-hpca-2018, ; cai2013error, ; cai2012flash, ; cai2017error, ; ha2015integrated, )) before or after its operations during the access of a read set (i.e., decompressing and preparing reads in the desired format). This is because (i) the duration of each SAGe process when accessing a read set, even if the read set occupies entire the SSD capacity, is significantly smaller than the SSD manufacturer-specified threshold for reliable retention age (e.g., one year (micron3dnandflyer, )), and because (ii) SAGe avoids read disturbance errors (cai2015read, ) by accessing a read set sequentially and with low reuse. By avoiding these management tasks during SAGe’s operations when preparing genomic data and leveraging SAGe’s efficient data placement, SAGe’s hardware accelerators can effectively leverage the SSD’s full bandwidth.
5.5. SAGe with Distributed Storage Systems
Due to SAGe’s efficient streaming access patterns and the ability to analyze different batches of genomic reads independently, SAGe is well-suited for integration with a distributed system. We elaborate on this by discussing three key aspects. The first aspect is the use of multiple SSDs. Since all data structures in SAGe are accessed sequentially, they can be evenly partitioned between multiple SSDs. This enables the efficient utilization of multiple SSDs in the system. The second aspect is inter-node communication. Since each batch of genomic reads can be analyzed independently, computation on the batches of data stored in each storage device is independent of data in other devices. This eliminates the need for frequent communication and synchronization between different nodes. The third aspect is accessing the storage system. From the perspective of the compute node, using distributed storage is the same as local storage (other than the latency/bandwidth properties across configurations). Therefore, SAGe can leverage the same type of interface command provisions (§5.3) for distributed storage systems as well.
6. Methodology
Evaluated Systems. We show the benefits of SAGe on end-to-end performance (preparing the data and performing read mapping on it) when it is integrated with a state-of-the-art read mapping accelerator, GEM (chen2023gem, ). For data preparation, we evaluate configurations with the following decompression tools: (ipigz: A parallel version (adler2015pigz, ) of gzip, a commonly-used general-purpose compressor; (ii(N)Spr: Spring (chandak2018spring, ) and NanoSpring (Meng2023, ), state-of-the-art compressors for short and long reads, respectively; (iii(N)SprAC: (N)Spr integrated with a BWT-accelerator. There are various accelerators available for BWT (e.g., (qiao2019fpga, ; guo2013gpu, ; zhao2017streaming, )). In our experiment, we consider an idealized accelerator that can fully eliminate the BWT execution time from (N)Spr; (iv0TimeDec: an ideal decompression tool (with zero time needed for decompression, and with a compression ratio similar to the state-of-the-art genomic compressor) outside the SSD, to show the limitations of decompression tools that cannot be efficiently implemented inside the SSD, regardless of their optimized performance outside the SSD; [^13] As a strongly conservative evaluation, 0TimeDec can also serve as an idealized representation of the closed-source industry software, called ORA (dragenora, ), recently developed for Illumina (illumina, ) short read (de)compression. ORA works similarly to other genomic compressors (as discussed in §2.3), but uses a different backend compressor for mismatch information, which still suffers from limitations for in-storage implementation (detailed in §4). We cannot directly use ORA as our short-read decompression baseline because it decompresses both DNA bases and quality scores, whereas SAGe, like many other tools, targets only the bases (as discussed in §5.1.5). Since ORA’s source code and implementation details are unavailable, modifying it to target only bases is infeasible, which prevents a fair quantitative comparison against it. (vSGSW: SAGe with its decompression implemented in software and running on the host system, to show the benefits of SAGe’s algorithmic optimizations; and (viSG: SAGe’s full implementation, with its decompression implemented in hardware. In all experiments except Fig. 13, SAGe’s hardware accelerator is inside the SSD. In Fig. 13, we compare both SAGe configurations (i.e., with its accelerator inside and outside the SSD) to evaluate their benefits on systems with different SSD bandwidth configurations. [^14] For fair comparisons, in all SAGe implementations, we use the same consensus sequence as (N)Spr. SAGe can flexibly work with any consensus sequence.
We demonstrate the benefits of SAGe’s in-storage implementation by evaluating its integration with a state-of-the-art ISP system for genomics, GenStore (mansouri2022genstore, ). GenStore prunes reads that do not require the costly read mapping process using low-cost and accurate in-storage filters, providing two key benefits: (i) mitigating data-movement overheads and (ii) reducing the computation burden from the rest of the system.
Datasets. We use real-world genomic data, with both short and long read sets. Storage systems must handle numerous read sets. Therefore, compressing each read set is essential to reduce the overall storage burden, regardless of their individual sizes. Thus, we analyze read sets of varying sizes, as listed in Table 3.
Performance. We design a simulator that models all of SAGe’s components, including accessing NAND flash chips, in-storage accelerators (both for SAGe’s hardware units and the in-storage filters (mansouri2022genstore, )), the SSD’s internal DRAM, and host-SSD and accelerator-SSD interfaces. Using this methodology, as also demonstrated in prior works (e.g., (mansouri2022genstore, ; park2022flash, ; megis, )), enables us to incorporate state-of-the-art system configurations in our analysis. We feed the latency and throughput of each component to this simulator. For the components in the hardware-based operations, we implement SAGe’s logic components in Verilog and synthesize them using the Synopsys Design Compiler (synopsysdc, ) with a 22 nm library (22gf, ). We use two state-of-the-art simulators, Ramulator (kim2016ramulator, ; ramulatorsource, ) to model the SSD’s internal DRAM, and MQSim (tavakkol2018mqsim, ; mqsimsource, ) to model the SSD’s internal operations. For the hardware read mapper, we use the throughput numbers reported by the original paper (chen2023gem, ). For the components in the software-based operations (software decompression), we measure performance on a real system, an ${}^{\text{\textregistered}}$ ${}^{\text{\textregistered}}$ 7742 CPU (amdepyc, ) with 128 physical cores and 1.5-TB DRAM. We measure the performance of the software tools on this high-end system using their best-performing thread counts.
SSD Configurations. In our real system experiments, we use a state-of-the-art SSD (samsungPM1735, ) with PCIe Gen4 (PCIE4, ). In Fig. 13, to analyze SAGe’s benefits in systems with SSDs of different interface bandwidths, we also use an SSD (samsung870evo, ) with a SATA3 interface (SATA, ). For the MQSim simulations of the in-storage tasks, we faithfully model the SSD configurations based on the configurations in Table 1.
Table 1. SSD configurations.
Specification PCIe SSD SATA SSD
General
48-WL-layer 3D TLC NAND flash-based SSD
4 TB capacity, 4 GB internal LPDDR4 DRAM (lpddr4, )
Bandwidth
4-lane PCIe Gen4 interface
7 GB/s sequential-read BW
1.2-GB/s channel I/O rate
SATA3 interface
560 MB/s sequential-read BW
1.2-GB/s channel I/O rate
NAND
Config
8 channels, 8 dies/channel,
4 planes/dies, 2,048 blocks/plane,
196 WLs/block, 16 KiB/page
8 channels, 8 dies/channel,
4 planes/dies, 2,048 blocks/plane,
196 WLs/block, 16 KiB/page
Latencies
Read (tR): 52.5 $\mu$s, Program (tPROG): 700 $\mu$s
Area and Power. For the accelerator logic units of SAGe, we use the area and power results from our Design Compiler synthesis. For the SSD, we use the power values of a Samsung 3D NAND flash-based SSD (samsung860pro, ). For DRAM, we use the power values of a DDR4 model (ddr4sheet, ; ghose2019demystifying, ). For the CPU cores, we use ${}^{\text{\textregistered}}$ µProf (microprof, ).
7. Evaluation
7.1. Performance
End-to-End Application Performance. Fig. 12 shows the end-to-end mapping performance (i.e., preparing data and performing read mapping on it). All configurations are integrated with a read mapping accelerator outside the SSD. For SAGe, we also evaluate an additional configuration in which it also integrates with a genomic in-storage filter (ISF). [^15] We use GenStore-EM for filtering exactly-matching short reads (mansouri2022genstore, ). We use GenStore-NM for filtering non-matching long reads in a contamination search read mapping use case presented in the original work (mansouri2022genstore, ). In this case, SAGe sends the unfiltered reads (i.e., reads that could not be pruned by ISF) in a decompressed 2-bit encoded format [^16] SAGe sends reads containing N in a 3-bit encoded format (§5.1.4). to the mapper outside the SSD. Speedup is normalized to the genomics-specific baseline (N)Spr.
Figure 12. End-to-end speedup for different read sets.
We make seven key observations. First, SAGe provides significant performance benefits compared to both general-purpose and genomics-specific baselines. SG leads to 12.3$\times$ and 3.9$\times$ average speedup over pigz and (N)Spr, respectively. Second, SG leads to 3.0$\times$ average speedup over (N)SprAC, while requiring only lightweight hardware and efficient access patterns. Third, by efficiently integrating with the in-storage filter, SG+ISF leads to 30.7$\times$, 9.9$\times$, and 7.5$\times$ average speedups compared to pigz, (N)Spr, and (N)SprAC, respectively. Fourth, due to its efficient sequential access patterns, SAGe’s implementation in software also leads to large speedups. On average, SGSW outperforms (N)Spr by 2.4$\times$. Fifth, despite its benefits, SGSW suffers from two key limitations: (i) decompression still significantly impacts end-to-end performance. SGSW leads to 1.6$\times$ average (up to 3.8$\times$) slowdown compared to the configuration without decompression overhead (0TimeDec), and (ii) by not executing inside the SSD, it does not enable integration with genomic ISP designs. SGSW performs, on average, 4.1$\times$ (up to 7.8$\times$) worse than SG+ISF. Sixth, SAGe can fully hide the decompression time in the end-to-end execution pipeline since the decompression time on a batch of reads in a read set is smaller than the mapping time for the previous batch. Note that, as explained in §3, I/O accesses, decompression, and read mapping work in a pipelined manner. Therefore, the end-to-end throughput is determined based on the slowest stage. 0TimeDec and SG achieving the same performance shows that in both cases, decompression is no longer the slowest stage of the end-to-end pipeline. The hardware implementation of SAGe decompression provides greater speedup than its software by handling bitwise operations (on SAGe’s arrays and guide arrays) more efficiently and managing data flow more effectively in a fine-grained manner. Seventh, regardless of how much decompression tools are optimized for performance outside the SSD, if they are unsuitable for in-SSD adoption (as discussed in §4), they miss out on the benefits of integration with genomic ISP systems. By efficiently integrating with ISF, SG+ISF achieves 2.5$\times$ larger average speedup compared to 0TimeDec.
Figure 13. Ablation analysis of SAGe’s optimizations.
Ablation Study. We perform an ablation study on SAGe’s optimizations by evaluating different configurations: (iSGSW: SAGe purely in software, i.e., only with its algorithmic optimizations, (iiSGout: SAGe with its hardware units, outside the SSD, (iiiSGin: SAGe with its hardware units, inside the SSD, (ivSGin+ISF: SGin integrated with the in-storage filter. [^17] As detailed in §6, we use GenStore (mansouri2022genstore, ) as the in-storage filter. Fig. 13 shows the end-to-end read mapping performance (i.e., preparing data with SAGe and performing read mapping on it). All configurations are integrated with a read mapper accelerator outside the SSD. Note that SAGe’s configurations outside the SSD only integrate with the outside-SSD mapper since their data preparation happens outside the SSD, preventing them from integration with ISF. We evaluate all configurations with a PCIe Gen4 SSD and a cost-optimized SATA SSD, which has significantly lower interface bandwidth than the PCIe SSD. Speedup is normalized to (N)Spr. We make four observations. First, due to its efficient access patterns, SGSW leads to large benefits compared to (N)Spr, providing 2.4$\times$ speedup. Second, SGout provides additional large speedups of up to 3.8$\times$ over SGSW. Third, for the case with low SSD bandwidth (i.e., the SATA SSD), SGout leads to larger benefits than SGin. Fourth, when integrated with ISF, SGin provides large benefits. SGin+ISF leads to better performance than SGout in all cases (providing up to 10.1$\times$ speedup), except when both of the following conditions are met: (i) the input and application do not largely take advantage of in-storage processing (i.e., in this case, ISF does not filter many reads in the read set), and (iithe SSD’s limited external bandwidth bottlenecks performance (e.g., RS1 and RS4 with the SATA SSD). In these cases, the SGout configuration should be used for better performance by decompressing data outside the SSD, thus avoiding moving larger decompressed data through the limited-bandwidth interface.
Performance with Multiple SSDs. Since all data structures in SAGe are accessed sequentially, they can be disjointly partitioned between multiple SSDs. Fig. 14 shows SAGe’s end-to-end performance with multiple SSDs, normalized to (N)Spr. First, SAGe maintains its large speedup when data is partitioned among multiple SSDs. Second, SAGe’s benefits improve for some datasets (RS3 and RS5) since more SSDs improve the performance of ISF, which is at the critical path of their end-to-end performance.
Figure 14. End-to-end speedup with different SSD counts.
Performance with Distributed Storage. Due to SAGe’s streaming accesses, the key factor of a distributed storage system impacting performance is its bandwidth. We consider two different configurations: (i) a high-end Lustre-based (lustre, ) distributed storage system connected to an InfiniBand network (infiniband, ), achieving a 10 GB/s bandwidth, and (ii) a low-end distributed storage system connected to an Ethernet network (10gbe_ethernet, ), achieving a 10 Gbps bandwidth. Fig. 15 shows the speedup of different SAGe configurations over (N)Spr. We observe that based on the available bandwidth, SGin or SGout can be selected to provide large benefits. For cases with limited bandwidth and with inputs that do not largely take advantage of in-storage processing (e.g., RS1 and RS4 with Ethernet-attached system), SGout should be used for better performance (2.6$\times$ average speedup over (N)Spr) by decompressing data outside the SSD. In all other cases, SGin should be selected to enable efficient integration with ISF, leading to 9.19$\times$ average speedup over (N)Spr.
Figure 15. Speedup with distributed storage configuration.
7.2. Area and Power
Table 2 lists the area and power consumption of SAGe’s accelerator units operating at 1 GHz. Although these units could operate at a higher frequency, their throughput is already sufficient because SAGe’s accelerator operations are bottlenecked by the NAND flash read throughput. SAGe’s accelerators consume a small area and power of 0.0023 mm2 and 0.95 mW at 22 nm node. The area overhead of SAGe hardware is only 0.7% of the three 28-nm ARM Cortex R4 cores (cortexr4, ) in a simple SSD controller (samsung860pro, ).  [^18] For a conservative comparison, we scaled SAGe’s area to 32nm node using (stillmaker2017Scaling, ).
Table 2. Area and power consumption of SAGe’s logic.
Logic unit # of instances Area [mm2] Power [mW]
Scan Unit 1 per channel 0.000045 0.014
Read Construction Unit 1 per channel 0.000017 0.080
Double Registers 1 per channel 0.00020 0.080
Control Unit 1 per channel 0.000029 0.025
Total for an 8-channel SSD - 0.002 0.95
7.3. Energy
We evaluate energy consumption for the end-to-end process of preparing data (reading data from the storage system, and decompressing and reformatting the read data) and performing read mapping on it. Read mapping is a fundamental process in the vast majority of genome analysis use cases (li2018minimap2, ; alser2020technology, ; cali2020genasm, ), for which it has been an optimization target in a large body of prior work (zhang2000greedy, ; slater2005automated, ; li2018minimap2, ; myers1999fast, ; marco2021fast, ; marcosola2023optimal, ; grootkoerkamp2024apa2, ). SAGe reduces the end-to-end energy by alleviating the energy overhead of preparing data for the energy-efficient read mapping accelerator. We evaluate the energy consumption based on the energy of the host processor and DRAM, accelerator units, SSD operations, and host/SSD/accelerator communication. Each part’s energy is based on its active/idle power and execution time. Fig. 16 shows the end-to-end energy reduction over pigz. We observe that, due to its lightweight structures and efficient access patterns, SG leads to a significant average energy reduction of 49.6$\times$, 24.6$\times$, and 18.8$\times$, over pigz, (N)Spr, and (N)SprAC, respectively.
Figure 16. End-to-end energy for different read sets.
7.4. Compression Ratio
Table 3 shows the compression ratios of pigz, (N)Spr, and SG. We observe that SG achieves (i) 2.9$\times$ better average compression ratio than pigz, and (ii) comparable compression ratios to the state-of-the-art genomic compression, (N)Spr, (with a modest 4.6% average reduction). SAGe achieves these high compression ratios while providing fast decompression speed. SAGe software (hardware) achieves 11.6$\times$ (53.0$\times$) and 3.3$\times$ (15.1$\times$) average decompression speedup over pigz  and (N)Spr, respectively. As shown in §7.1, this faster decompression leads to substantial improvement in the end-to-end throughput of genome analysis pipelines.
Table 3. Compression ratios for different read sets.
Label Read Set (Short, Long) Uncompressed Size (MB) Compression Ratio
PigZ (adler2015pigz, ) (Nano)Spring (Meng2023, ) SAGe
RS1 SRR870667_2 (Motamayor2013, ) (S) 5 000 3.39$\times$ 24.8$\times$ 22.8$\times$
RS2 ERR194146_1 (eberle2017reference, ) (S) 79 000 12.5$\times$ 40.2$\times$ 36.8$\times$
RS3 SRR2052419_1 (zook2016extensive, ) (S) 4 000 3.41$\times$ 7.2$\times$ 7.1$\times$
RS4 PAO89685_sampled (ontopendata, ) (L) 12 000 3.93$\times$ 4.8$\times$ 4.5$\times$
RS5 ERR5455028 (Belser2021, ) (L) 88 400 3.5$\times$ 7.6$\times$ 7.8$\times$
To show the impact of different optimizations, Fig. 17 presents the size breakdown of the reads’ mismatch information [^19] Rev refers to a bit marking if a read matches in reverse to the consensus sequence. for a short (RS2) and long read set (RS4) in five settings: (iNO, with no optimization on the raw mismatch information; (iiO1, with matching position optimization (§5.1.3) added; (iiiO2, with mismatch positions and count optimizations (§5.1.1) added; (ivO3, with mismatching base and type optimizations (§5.1.2) added; and (vO4, with corner case optimizations (§5.1.4) added.
Figure 17. Impact of different SAGe optimizations.
8. Related Work
Accelerating Genome Analysis. Many prior works accelerate genome analysis via algorithmic optimizations (e.g., (zhang2000greedy, ; slater2005automated, ; li2018minimap2, ; myers1999fast, ; marco2021fast, ; marcosola2023optimal, ; grootkoerkamp2024apa2, )), or hardware acceleration using GPUs (e.g., (jia2011metabing, ; kobus2021metacache, ; wang2023gpmeta, ; kobus2017accelerating, ; Su2012, ; su2013gpumetastorms, ; Yano2014, ; Gamaarachchi2020_f5c, ; cheng2018bitmapper2, ; houtgast2018hardware, ; houtgast2017efficient, ; zeni2020logan, ; ahmed2019gasal2, ; nishimura2017accelerating, ; de2016cudalign, ; liu2015gswabe, ; liu2013cudasw++, ; wilton2015arioc, ; liu2009cudasw++, ; liu2010cudasw++, )), FPGAs (e.g., (saavedra2020mining, ; zhang2023genomix, ; cervi2022metagenomic, ; shih_haru_2023, ; fujiki2020seedex, ; banerjee2018asap, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, ; liyanage2023efficient, )), ASICs (e.g., (turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, )), processing-in-memory (e.g., (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, )), and ISP (e.g., (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )). There are also many accelerators for specific parts of the genome analysis pipeline  (saavedra2020mining, ; markus2020benchmarking, ; subramaniyan2021accelerated, ; fujiki2020seedex, ; huangfu2018radar, ; khatamifard2021genvom, ; cali2020genasm, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, ; cheng2018bitmapper2, ; houtgast2018hardware, ; houtgast2017efficient, ; zeni2020logan, ; ahmed2019gasal2, ; nishimura2017accelerating, ; de2016cudalign, ; liu2015gswabe, ; liu2013cudasw++, ; liu2009cudasw++, ; liu2010cudasw++, ; wilton2015arioc, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fujiki2020seedex, ; banerjee2018asap, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, ). SAGe is orthogonal to these works and can be flexibly integrated with them to further improve their benefits.
Genomics-Specific Compression. Due to inefficiencies of general-purpose compression for genomic data, a large body of works (e.g., (chandak2018spring, ; Deorowicz2020, ; lan2021genozip, ; alyami2019lfastqc, ; kowalski2019pgrc, ; roguski2018fastore, ; chandak2017compression, ; cogo2021genodedup, ; Meng2023, ; kokot2022colord, ; dufort2020enano, ; dufort2021renano, ; karasikov2022lossless, ; vandamme2024tinted, ; dragenora, )) improves compression ratios via genomics-specific compression. As detailed in §4, some works (e.g., (guo2013gpu, ; qiao2019fpga, ; zhao2017streaming, ; jiang2021exma, ; arram2015fpga, ; wang2018accelerating, ; chen2023efficient, ; leavline2013hardware, )) accelerate certain computational kernels that are widely used in state-of-the-art genomic compressors. Despite their benefits, these works have two key limitations. First, as detailed in §4, they are unsuitable for implementation in the resource-constraint environment inside modern SSDs. Second, they do not fully alleviate end-to-end data preparation overhead for genomic accelerators (detailed in §4 and evaluated in §7). While some works (rajarajeswari2011dnabit, ; saada2016dna, ) use algorithms that do not rely on expensive units, they achieve poor compression ratios (§4 ).
General-Purpose Compression. Many works propose general-purpose compression in software (e.g., (collet2018zstandard, ; pavlov20167, ; Brotli, ; oswal2016deflate, ; goyal2021dzip, ; goyal2018deepzip, ; chen2024ha, )) or hardware (e.g., (bartik2015lz4, ; liu2018data, ; fowers2015scalable, ; chen2021fpga, ; angerd2022gbdi, ; gao2024beezip, ; karandikar2023cdpu, ; 9499902, ; abali2020data, )). However, they achieve poor compression ratios for genomic data compared to genomic compressors (hernaez2019genomic, ; zhu2013high, ). For example, the compression accelerators Intel QAT (intelqat, ) and IBM zEDC (ibmzedc, ) lead to 2.6×2.6$\times$2.6 × and 2.7×2.7$\times$2.7 × lower average compression ratios, respectively, than genomic compressors on our evaluated datasets. Due to their limitations in leveraging longer-range similarities common in DNA data (hernaez2019genomic, ; zhu2013high, ), even state-of-the-art strong general-purpose compression algorithms do not achieve the compression ratios on par with genomic compression tools. For example, we evaluate xz (xz, ) (a state-of-the-art LZMA-based compression tool) and zstd (collet2018zstandard, ) (a recently-developed compression algorithm), both with their highest compression levels, and show that they achieve 2.13$\times$ lower average compression ratios than the state-of-the-art genomic compressors across our genomic datasets. By leveraging the unique features of genomic data, SAGe achieves compression ratios similar to the state-of-the-art genomic compression (with a modest reduction of 4.7%, as shown in §7.4), while also providing faster decompression. Our evaluations show that the software implementation of SAGe achieves 2.1$\times$ and 3.9$\times$ average speedup over xz and zstd, respectively.
In-Storage Processing. Many prior works propose ISP architectures for various applications (liang2019cognitive, ; kim2020reducing, ; lim2021lsm, ; li2021glist, ; wang2016ssd, ; lee2020neuromorphic, ; kang2021s, ; han2021flash, ; wang2022memcore, ; wang2018three, ; han2019novel, ; choi2020flash, ; mailthody2019deepstore, ; pei2019registor, ; jun2018grafboost, ; do2013query, ; seshadri2014willow, ; kim2016storage, ; riedel2001active, ; riedel1998active, ; lee2022smartsage, ; jeong2019react, ; jun2016storage, ; li2023ecssd, ; wang2024beacongnn, ; jang2024smart, ; lee2024presto, ; mahapatra2024instoragedomainspecificaccelerationserverless, ; 10.1145/3689336, ; 10.1145/3632951, ; 10.1145/3631882.3631896, ), such as machine learning (e.g., (li2023ecssd, ; liang2019ins, ; lee2022smartsage, ; wang2024beacongnn, ; jang2024smart, ; pan2024instinferinstorageattentionoffloading, )), genome analysis (e.g., (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )), and graph analysis (e.g., (jun2018grafboost, )). Some works integrate additional processing units into the SSD controller (gu2016biscuit, ; kang2013enabeling, ; wang-eurosys-2019, ; acharya-asplos-1998, ; keeton-sigmod-1998, ; mailthody2019deepstore, ; pei2019registor, ; jun2018grafboost, ; do2013query, ; seshadri2014willow, ; kim2016storage, ), utilize flash cells (park2022flash, ; choi2020flash, ; han2019novel, ; shim2022gp3d, ; merrikh2017high, ; tseng2020memory, ; wang2018three, ; lue2019optimal, ; parabit_2021_gao, ; yu2024cambriconllmchipletbasedhybridarchitecture, ; ares2024, ), or tightly integrate SSDs with FPGAs (koo-micro-2017, ; jun2015bluedbm, ; torabzadehkashi2019catalina, ; lee2020smartssd, ; ajdari2019cidr, ; 10.1145/3649329.3658265, ) or GPUs (cho2013xsd, ). SAGe is orthogonal to prior ISP genomics works and enables their adoption without relying on large data to be stored uncompressed.
9. Conclusion
We propose SAGe, an algorithm-system co-design for capacity-efficient storage and high-performance access of large-scale genomic data. We leverage the features of genomic data in compression mechanism and storage architecture design. SAGe can be seamlessly integrated with a broad range of genome analysis systems. SAGe significantly improves the end-to-end performance and energy efficiency of genome analysis accelerators, while achieving high compression ratios and at a low hardware cost.